find the divergence of of the field given by a=1/r in spherical coordinate q


1
Expert's answer
2021-05-17T19:01:23-0400

"\nabla\vec A=\frac{1}{r^2}\frac{\partial(r^2A_r)}{\partial r}+\frac{1}{r\sin\theta}\frac{\partial}{\partial\theta}(A_\theta\sin\theta)+\frac{1}{r\sin\theta}\frac{\partial A_\varphi}{\partial\varphi}
=\frac{1}{r}"